Analyze the free-body diagrams (FBDs) and place them in the indicated order according to the magnitudes of their net forces.AF, (N) у 60 60 50 50 Ả 40 40 ſ K 30 30 20 20 { 10- 1 F (N) 10- F (N) -50 -40 -30 -20 -10 10 20 30 40 50 10 20 30 40 50 -50 -FBD 3 FBD 4 Smallest net force FBD 3 FBD 4 FBD 1 FBD 2 Largest net force Answer Bank
Answer #1

FBD1 FBD2 Ā = -301 +40 B = 20 e = 20 1 +105 D = - 50 Ř = + B tct = -107 + 205 lel = vion too 22.36 N Ē=-yoi F = 40 it 50 G =